# Panoramic Out-of-Distribution Segmentation

### v1

@@ 5.6 Ablation Study
Each Component in POS. Table V presents the component-wise ablation results of the technical novelties included in POS. Using the image encoder alone as the backbone, the model demonstrates strong potential for omnidirectional visual perception. However, it still struggles with semantic ambiguity and background interference in complex panoramic scenes. By incorporating Prompt-based Restoration Attention (PRA), the model significantly improves pixel-level semantic associations through prompt distribution learning, enhancing panoramic perception. This results in an increase in AuPRC to $83.94\%$ ($+7.06\%$) and an improvement in mIoU to $62.50\%$ ($+0.77\%$), highlighting the effectiveness in mitigating dense pixel semantic blurring. Further incorporation of Bilevel Prompt Distribution Learning (BPDL) leads to a synergistic optimization: PRA enhances inter-class discrimination through semantic decoupling, while BPDL preserves topological consistency via manifold learning. This combined mechanism enables the model to achieve groundbreaking performance in the PanOoS task: $85.56\%$ AuPRC, $0.45\%$ FPR95, and $63.30\%$ mIoU. Not only does it overcome the coupling challenges of semantic ambiguity and geometric distortion that traditional pinhole-OoS methods cannot deal with, but it also establishes a new benchmark for panoramic scene understanding.
@@
We further empirically evaluate the impact of fine-tuning parameters during the optimization of the prompt distribution space. As shown in Table VIII(a), performance evaluations on the DenseOoS dataset for PanOoS demonstrate that the model reaches optimal performance after approximately $30k$ iterations, after which overfitting begins. For the outlier data exposure control parameter $p_{out}$, experiments reveal that when $p_{out}{>}0.3$, both closed-set performance and out-of-distribution segmentation performance deteriorate. Based on these results, we selected $0.3$ as the optimal value, balancing outlier detection with closed-set performance.
Moreover, a comparison of fine-tuning strategies (see Table VIII(c)) shows that the panoramic-oriented disentanglement (Pixel Dec.) outperforms other strategies. Specifically, fine-tuning only the MLP and linear layers (w/o) or the Transformer decoder (Trans. Dec.) fails to effectively address the complexities of panoramic images. While additional fine-tuning of all decoders performs comparably in FPR95, it does not achieve optimal results on AuPRC and mIoU, decreasing segmentation accuracy. This suggests that excessive fine-tuning can introduce semantic ambiguity, compromising segmentation accuracy. These findings verify the effectiveness of our proposed fine-tuning strategy in preserving semantic consistency and enhancing out-of-distribution detection.
Features and layers in Transformer Decoder. To further analyze the most favorable settings for PanOoS, we conduct additional ablation experiments on the decoded features and corresponding number of layers in Transformer decoder. As shown in Table IX, the results align with previous work: by decoding only semantically rich features and reducing the number of decoder layers, semantic confusion can be effectively alleviated, which helps to better align the logit scores and results in better outlier performance [24]. Specialized object queries perform worse with more decoder layers. This also supports our view that, while ensuring semantic clarity, enhancing the model’s perception of panoramic pixels can significantly improve the performance of PanOoS.
Features Layer AuPRC FPR95 mIoU
$f_{4}$ 1 85.56 0.45 63.30
3 71.18 0.68 57.72
6 76.46 0.64 57.33
$f_{2}$, $f_{3}$, $f_{4}$ 3 72.38 0.94 58.96
6 63.60 1.01 58.66
9 65,91 1.06 58.79
Table IX: Comparison of different features in Transformer Decoder. The $f_{2}$ and $f_{3}$ are feature maps with stride $8(c_{8})$ and $16(c_{16})$ processed by DeformAttn layers, respectively.
6 Conclusion
@@
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. U21A20518, No. U23A20341, and No. 62473139.
@@ Appendix B Benchmarks and Experiments
B.1 More Details of Datasets
As shown in Table B.1, we present the specific types of outliers in DenseOoS. DenseOoS is based on the designs in POC [80] and additionally includes $5$ other common anomalous objects (cattle, duck, cock, box, and rock). In addition, we also list some types of outliers in QuadOoS. As present in Fig. B.1, the outliers mainly consist of two parts: human-defined and existing in the neighborhood.
@@ Appendix C More Quantitative Results
C.1 More in-distribution segmentation results
@@
To provide a thorough analysis of the panoramic segmentation and demonstrate that excellent closed-set segmentation performance ensures effective out-of-distribution segmentation, we present the accuracy results for each inlier class in the DenseOoS dataset, as shown in Table C.1. To benchmark our model, we conduct a comparative study with the previous state-of-the-art panoramic semantic segmentation methods, namely DAFormer [31], Trans4PASS [8], Trans4PASS+ [32], and 360SFUDA++ [46]. DAFormer introduces domain-adaptive transformers to improve semantic segmentation across domains. Trans4PASS and its enhanced variant, Trans4PASS+, employ transformer-based architectures specifically designed to mitigate distortions and exploit contextual dependencies in panorama. 360SFUDA++ represents an advanced source-free domain adaptation framework based on self-training, tailored to panoramic scenarios. While these methods achieve good performance under closed-set conditions, they experience notable degradation in the presence of out-of-distribution regions, owing to their limited generalization beyond the predefined semantic space. This limitation underscores the importance of enhancing robustness to unexpected content in real-world panoramic scenes.

### v2

@@ V-F Ablation Study
Each Component in POS. Table V presents the component-wise ablation results of the technical novelties included in POS. Using the image encoder alone as the backbone, the model demonstrates strong potential for omnidirectional visual perception. However, it still struggles with semantic ambiguity and background interference in complex panoramic scenes. By incorporating Prompt-based Restoration Attention (PRA), the model significantly improves pixel-level semantic associations through prompt distribution learning, enhancing panoramic perception. This results in an increase in AuPRC to $83.94\%$ ($7.06\%$) and an improvement in mIoU to $62.50\%$ ($0.77\%$), highlighting the effectiveness in mitigating dense pixel semantic blurring. Further incorporation of Bilevel Prompt Distribution Learning (BPDL) leads to a synergistic optimization: PRA enhances inter-class discrimination through semantic decoupling, while BPDL preserves topological consistency via manifold learning. This combined mechanism enables the model to achieve groundbreaking performance in the PanOoS task: $85.56\%$ AuPRC, $0.45\%$ FPR95, and $63.30\%$ mIoU. Not only does it overcome the coupling challenges of semantic ambiguity and geometric distortion that traditional pinhole-OoS methods cannot deal with, but it also establishes a new benchmark for panoramic scene understanding.
@@
Bilevel Prompt Distribution Learning. The proposed BPDL belongs to the domain of metric learning. Unlike traditional methods that rely on manually defined or learned centroid embeddings as class prototypes, BPDL uses semantically rich text embeddings as prototypes. This approach allows for a coarse-grained predefinition of distribution regions for each class. Moreover, BPDL utilizes a bilevel learning framework for distribution optimization: the lower-level optimization refines the inter-class distributions for known classes, while the upper-level optimization adjusts the global distribution space. As shown in Table VII, each component of the $\mathcal{L}_{bpdl}$ contributes to spatial distribution optimization, with the $\mathcal{L}_{ind}$ and $\mathcal{L}_{pixel}$ loss terms leading to particularly notable performance gains (AuPRC: $2.95\%/{+}2.48\%$, mIoU: $0.63\%/{+}0.41\%$). Ablation experiments confirm that integrating all loss components results in optimal performance across all evaluation metrics.
((a))
((b))
((c))
((d))
@@
Outlier Exposure Fine-tuning. We further empirically evaluate the impact of fine-tuning parameters during the optimization of the prompt distribution space. As shown in Table LABEL:tab:ablationIter, performance evaluations on the DenseOoS dataset for PanOoS demonstrate that the model reaches optimal performance after approximately $30k$ iterations, after which overfitting begins. For the outlier data exposure control parameter $p_{out}$, experiments reveal that when $p_{out}{>}0.3$, both closed-set performance and out-of-distribution segmentation performance deteriorate. Based on these results, we selected $0.3$ as the optimal value, balancing outlier detection with closed-set performance.
Moreover, a comparison of fine-tuning strategies (see Table LABEL:tab:ablationStr) shows that the panoramic-oriented disentanglement (Pixel Dec.) outperforms other strategies. Specifically, fine-tuning only the MLP and linear layers (w/o) or the Transformer decoder (Trans. Dec.) fails to effectively address the complexities of panoramic images. While additional fine-tuning of all decoders performs comparably in FPR95, it does not achieve optimal results on AuPRC and mIoU, decreasing segmentation accuracy. This suggests that excessive fine-tuning can introduce semantic ambiguity, compromising segmentation accuracy. These findings verify the effectiveness of our proposed fine-tuning strategy in preserving semantic consistency and enhancing out-of-distribution detection.
VI Conclusion
@@ Appendix B Benchmarks and Experiments
B-A More Details of Datasets
B-A1 DenseOoS
As shown in Table B.1, we present the specific types of outliers in DenseOoS. DenseOoS is based on the designs in POC [81] and additionally includes $5$ other common anomalous objects (cattle, duck, cock, box, and rock). To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$ independent screening $\rightarrow$ voting” workflow.
The overall process of dataset generation and construction is as follows. First, all text prompts for image inpainting (a good photo of a $\{class\}$) are determined based on the DensePASS validation dataset; Next, we select appropriate regions to be inserted (e.g., the road and sidewalk) and the selected rectangular area is scaled to match the resolution of the real panoramic images. Then, we set different random seeds and use the POC algorithm to batch-generate images and outlier annotations. Finally, the suitable images for constructing the DenseOoS dataset are determined through preliminary manual screening, independent screening, and majority voting.
B-A2 QuadOoS
In addition, we also list some types of outliers in QuadOoS. As present in Fig. 16, the outliers mainly consist of two parts: human-defined and existing in the neighborhood.
@@ Appendix C More Quantitative Results
C-A More in-distribution segmentation results
@@
To provide a thorough analysis of the panoramic segmentation and demonstrate that excellent closed-set segmentation performance ensures effective out-of-distribution segmentation, we present the accuracy results for each inlier class in the DenseOoS dataset, as shown in Table C.1. To benchmark our model, we conduct a comparative study with the previous state-of-the-art panoramic semantic segmentation methods, namely DAFormer [34], Trans4PASS [12], Trans4PASS+ [2], and 360SFUDA++ [48].
DAFormer introduces domain-adaptive transformers to improve semantic segmentation across domains. Trans4PASS and its enhanced variant, Trans4PASS+, employ transformer-based architectures specifically designed to mitigate distortions and exploit contextual dependencies in panorama. 360SFUDA++ represents an advanced source-free domain adaptation framework based on self-training, tailored to panoramic scenarios. While these methods achieve good performance under closed-set conditions, they experience notable degradation in the presence of out-of-distribution regions, owing to their limited generalization beyond the predefined semantic space. This limitation underscores the importance of enhancing robustness to unexpected content in real-world panoramic scenes.
@@
C-B Additional Ablation Study
Features Layer AuPRC FPR95 mIoU
$f_{4}$ 1 85.56 0.45 63.30
3 71.18 0.68 57.72
6 76.46 0.64 57.33
$f_{2}$, $f_{3}$, $f_{4}$ 3 72.38 0.94 58.96
6 63.60 1.01 58.66
9 65.91 1.06 58.79
TABLE C.2: Comparison of different features in Transformer Decoder. The $f_{2}$ and $f_{3}$ are feature maps with stride $8(c_{8})$ and $16(c_{16})$ processed by DeformAttn layers, respectively.
Features and layers in Transformer Decoder. To further analyze the most favorable settings for PanOoS, we conduct additional ablation experiments on the decoded features and the corresponding number of layers in the transformer decoder. As shown in Table C.2, the results align with previous work: by decoding only semantically rich features and reducing the number of decoder layers, semantic confusion can be effectively alleviated, which helps to better align the logit scores and results in better outlier performance [27]. Specialized object queries perform worse with more decoder layers. This also supports our view that, while ensuring semantic clarity, enhancing the model’s perception of panoramic pixels can significantly improve the performance of PanOoS.